OpenStudy (anonymous):

Tonya is twice Kevin's age in 3 years, tonya will be 17 . How old is kevin

OpenStudy (anonymous):

Let tonya = T Let Kevin = K 1) T = 2K (Tonya is twice Kevin) 2) T+3 = 17 (Tonya's age + 3 = 17) T = 17-3 = 14. Now from 1), T = 2K. We know T = 14, so: 14 = 2K K = 14/2 = 7 C. 7. is the answer

OpenStudy (anonymous):

hope this helps